# Floquet-engineered Emergent Massive Nambu–Goldstone Modes

### v3

@@ 4 Different Scaling of Linewidth
In the main text, Fig. 4, we find that the linewidth of DSF scales as $T^{3}$, and we give one possible explanation using a hydrodynamic theory. One key assumption we made is that the attenuation $D_{0}$, induced by explicit SB, is dominated by $Q^{(2)}$. Together with the scaling of the gap size, one estimates the damping rate as $\Gamma\sim T^{3}$.
However, this scaling may not be universal, and it can depend on the driving parameters and the randomness in the initial state ensemble.
@@
Figure S3: The width $\Gamma$ exhibits a power-law scaling in the high frequency regime; the dashed line with has fitted scaling $\alpha\approx 4$. We use $N{=}1000$, $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}-2$ for numerical simulations. Ensemble averaging over $100$ realizations is performed with $\theta_{j}{=}0.1\pi$ and azmuthal angles $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@ 5 Thermalization and Hierarchical Symmetry Breaking in Prethermal Plateaus
Floquet systems generally heat up to infinite temperature, due to the absence of energy conservation. However, this can take a long time, with a timescale that normally increases exponentially with the driving frequency [63, 64]. In Fig. 1 in the main text, to show the persistent periodic oscillation of the magnetization we use a small $T$, which indeed significantly delays the onset of heating. During the time window that we can numerically simulate, the system does not heat up to infinite temperature, but stays in a prethermal plateau. Here, instead, we use a smaller driving frequency and stronger symmetry breaking strength $J_{x}$ to speed up heating. As shown in Fig. S4, after the initial transient oscillation, the system relaxes towards a prethermal plateau. Eventually, magnetization in three directions decays to zero, confirming the eventual heat death.
Figure S4: Dynaimics of the 1D driven Heisenberg-type model, Eq. (3), that shows the eventual thermalization. We use $N{=}1000$, $T{=}2.4$ $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}-2$, $J_{x}{=}5$ for numerical simulations. Ensemble averaging over $100$ realizations is performed with $\theta_{j}{=}0.1\pi$ and azmuthal angles $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@
$U_{F}=(e^{iH_{0}\frac{T}{6}}e^{iH_{1}\frac{T}{6}})e^{-iH_{2}\frac{T}{6}}(e^{-iH_{0}\frac{T}{6}}e^{-iH_{1}\frac{T}{6}})e^{-iH_{2}\frac{T}{6}},$ (S.12)
@@
$H_{2}=-J_{2}\sum_{j}\bm{S}_{j}\cdot\bm{S}_{j+1},\ \ H_{1}=-H_{2}+J_{1}\sum_{j\in even}S_{j}^{z}S_{j+2}^{z},\ \ H_{0}=J_{x}\sum_{j\in odd}S_{j}^{x}S_{j+2}^{x}.$ (S.13)
Figure S5: (a) Dynamics of the Floquet system, Eq. (S.12). The prethermal plateaus for magnetization in $z$ direction survives longer than the other two, showing a hierarchy in symmetry breaking phenomenon. (b) We define the lifetime $\tau$ as the time when the magnetizations drop below the threshold value $0.16$. $\tau$ exhibits a power law dependence on the driving period, following the form $\tau\sim T^{-\alpha}$ where $\alpha\approx 2$ for the $x$ (green dots) and $y$ (not shown but nearly coinciding with the x-component) components, and $\alpha\approx 4$ for the $z$ (orange dots) components. This scaling is consistent with the Fermi’s Golden Rule expectation. We use $N{=}100$, $J_{2}{=}1$, $J_{1}{=}1$, $J_{x}{=}1$ for numerical simulations. Ensemble averaging over $800$ realizations is performed with the initial state where $40\%$ of the sites are prepared as $S_{j}^{x}=S_{j}^{y}=S_{j}^{z}=1/\sqrt{3}$, while the remaining $60\%$ are randomly sampled.
Fig. S5 (a) shows the numerical time evolution of the total magnetization. From a sufficiently random initial ensemble, all components of the magnetization quickly settle into prethermal plateaus before eventually heating up to infinite temperature. A hierarchy in the prethermal lifetime is observed, and the $z$-component exhibits a more stable plateau that persists longer than those of the $x$ and $y$ components. Moreover, Ref. [56] predicts that the prethermal lifetime $\tau$ can be parametrically prolonged by using smaller driving period $T$, $\tau_{x/y}\sim T^{-2}$ and $\tau_{z}\sim T^{-4}$, following the Fermi Golden Rule argument. There, this scaling has been numerically verified in quantum systems of a small size. In Fig. S5 (b) we also observe this scaling law in our classical systems where finite-size effects are negligible.
@@ 6 Classical System with Three-body Interactions
To realize emergent mNGs, explicit SB processes in $Q^{(1)}$ are not limited to being a generator of the non-Abelian group $G_{2}$, as long as $Q^{(1)}$ still preserves a subsymmetry of $G_{2}$. Here, we present an explicit example that realizes the symmetry-breaking process, $O(3)-O(2)-E$, and $Q^{(1)}$ involves three-body interactions which is not a simple generator of the $O(3)$ group. We still observe emergent mNGs with controllable dispersion relation that can be tuned by $T$. However, those mNGs have a shorter lifetime that scales linearly with the driving frequency.
@@
We obtain the effective Hamiltonian up to order $\mathcal{O}(T)$
@@
$Q^{(0)}$ preserves $O(3)$ and $Q^{(1)}$ contains only the exchange terms and the $z$-component terms, therefore reducing the symmetry to $O(2)$ around the $z$-axis. We take $J_{2}>0$ such that the ground state features the FM order along the positive (negative) $z$-direction if $J_{1}$ and $h_{y}$ have the same (opposite) sign. The EOM generated by the truncated Hamiltonian Eq. (S.15) read as
@@
$\omega=\pm\left[\frac{2J_{2}}{5}(1-\cos{qa})+\frac{J_{1}h_{y}T}{200}(3+2\cos{qa}+2\cos{2qa})\right],$ (S.18)
A quadratic dispersion appears for modes of long wavelength when $q\rightarrow 0$ and a gap is now opened at $q=0$ of size $|\frac{7J_{1}h_{y}T}{200}|$ that is linear in $T$.
@@
Figure S6: Dynamics of the total magnetization in a 1D driven three-body interaction model, cf. Eq. (S.14). Magnetization in the $(x,y)$-plane (red and blue) rotates periodically, confirming the existence of mNGs. $\langle S_{z}^{\text{tot}}\rangle/N$ (green) oscillates on top of the conserved value (dashed line) due to higher order effects. We use a system of size $N=1000$ and parameters $T{=}1$, $J_{2}{=}1$, $h_{y}{=}1$, $J_{1}{=}1$, $J_{x}{=}1$. We perform ensemble averages over $100$ realizations where initial spins deviate from the $z$-axis by one small angle $\theta_{j}=0.05\pi$, and their azimuthal angles $\phi_{j}$ are randomly sampled in the range $(-\pi/3,\pi/3)$.
@@
For even longer times, the damping of mNG modes broadens the spectrum and induces a finite linewidth $\Gamma$ in DSF. We extract $\Gamma$ for $q=0$ by performing Fourier transform on the dynamics of the total magnetization components over time, and plot it in Fig. S9 for varying $T$. The linewidth is determined by averaging the full-width at $10\%$, $15\%$, and $20\%$ of the maximum intensity of the Fourier peak, with their standard deviation serving as an error bar. $\Gamma$ exhibits a power-law scaling $T^{\alpha}$. However, in contrast to Fig. 4, here the scaling exponent is about 1, cf. Fig. S9, and hence mNGs are not very robust. Such a scaling implies that a $Q^{(1)}$ of order $\mathcal{O}(T)$ can destablize mNGs, even if it preserves the $O(2)$ symmetry. The approximation that one can treat the three-body interaction as an effective $O(3)$ generator fails after a time scale that scales as $\mathcal{O}(T^{-1})$. Therefore, the damping of mNGs becomes notable shortly after the failure of linearization. We simulate the spin dynamics by using the truncated Hamiltonian, Eq. (S.15), and the results (green dots) are depicted in Fig. S9(a). In the high frequency regime, the linewidth for this quenched protocol closely mimics the Floquet results, confirming that $Q^{(1)}$ indeed dominates the damping of mNGs. For slightly larger drive period, the Floquet protocol generates a smaller linewidth, i.e., a longer lifetime. It essentially implies that $Q^{(2)}$ indeed stabilizes these mNGs, an counter intuitive observation since higher order terms normally speed up heating and hence shorten the lifetime of quasi-particles.
Figure S9: 1D driven three-body interaction model, cf. Eq. (S.14): (a) The linewidth $\Gamma$ exhibits a power-law scaling in the high-frequency regime with the exponent closing to $1$. The results of the quenched protocol closely mimic the Floquet protocol in the high-frequency regime. Floquet protocol generates a smaller linewidth for large $T$. $\theta_{j}{=}0.1\pi$. (b) The monotonic dependence of $\lambda$ on $\theta$. The ratio $\lambda$ can be suppressed by applying smaller spatial randomness. Various $\theta_{j}=\theta$ are used for characterizing the spatial randomness. We use $N{=}1000$, $T{=}1$, $J_{2}{=}1$, $h_{x}{=}1$, $J_{1}{=}1$ and $J_{x}{=}1$. $100$ ensemble averages are performed with azimuthal angles $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@ 7 Ferromagnetic Quantum Spin Model
In the main text, we use a classical spin model to verify the existence of mNGs. However, the Floquet protocol equally applies to both quantum and classical systems. Therefore, we expect similar dynamical phenomena in the quantum system for the FM case. We verify the existence of magnon excitations by applying the protocol Eq. (1) to a quantum many-body system of spin-$\frac{1}{2}$, with the Hamiltonian
@@
where we use Pauli operators.
We use Quspin [89] for numerical simulation and the system initially has all spins pointing up. Then we perform a single-site rotation on all spins around the $x$-axis by $e^{-i\sum_{j}\sigma_{j}^{x}\theta_{j}/2}$, where the angle $\theta_{j}$ is chosen in a small range, yielding a sufficiently large magnetization density in $z$-direction and non-vanishing initial magnetization in the $(x,y)$-plane. The spatial randomness can mimic the thermal fluctuation that may exist in real quantum simulator platforms.
Figure S10: (a) Dynamics of the total magnetization of the quantum spin chain, Eq. (S.19), with $T=2.4$. Magnetization in the $(x,y)$-plane (red and blue) rotates at early times, and eventually decays. (b) Dynamics of the total magnetization with a smaller driving period $T=1.2$. Persistent periodic rotation of magnetization in the $(x,y)$-plane (red and blue) confirms the existence of mNGs. (c) Linear dependence of the gap size versus $T$. We use a system of size $N=20$ and parameters $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}-2$, $J_{x}{=}1$. We perform ensemble averages over $100$ realizations where all spins rotate around the $x$-axis by $e^{-i\sum_{j}\sigma_{j}^{x}\frac{\theta_{j}}{2}}$ with $\theta_{j}$ randomly sampled within the range $(-\pi/100,\pi/100)$.
@@
We have provided a comprehensive analysis of FM in the main text, categorized as case 1. For case 2, we investigate the properties of the mNG modes in the Floquet system with AFM interaction, $J_{2}<0$. Significant distinctions between the two cases can be observed. For AFM, if we consider only the first two orders of the effective Hamiltonian, there exists a gapless Goldstone mode with linear dispersion and a gapped mNG with quadratic dispersion  [66]. Taking higher-order terms into account, the gapless mode disappears and the spectrum of the gapped mode is slightly modified. The gap linearly depends on the drive period $T$ in the high-frequency regime. Furthermore, we note that the stability of the gapped mode is quite different from the FM case.
@@ 1 Spin-wave Theory
where $J<0$ for AFM interaction. If $h_{z}=0$, there is one NG mode with linear dispersion. For non-zero $h_{z}$ the ground state of the system features all spins having the same $z$-component with $S_{j}^{z}=\frac{h_{z}}{4J}$ and being antiparallel in $(x,y)$-plane for $|h_{z}|<-4J$. For $|h_{z}|>-4J$, the ground state state is just the FM state along the $z$-axis. We focus on the former situation, since in our Floquet system the effective $z$-field strength linearly scales with $T$, which is a perturbatively small quantity. The Hamiltonian leads to the EOM
@@ 2 Numerical Verification of Massive Goldstone Modes
Figure S11: Dynamical structure factor (DSF) through the Brillouin zone of AFM model. (a) Obtained from spin-dynamics simulation of a time-independent Hamiltonian, Eq. (S.20), with parameters $J{=}{-}1/5$, $h_{z}{=}1/25$. A gapped mode with quadratic dispersion and a gapless mode with linear dispersion appear in accordance with the theoretical prediction. (b) Floquet spin-dynamics simulation with $T{=}2$, $J_{2}{=}{-}1$, $h_{x}{=}2$, $h_{y}{=}2$ and $J_{x}{=}1$, which gives the same parameters as (a). However, this spectrum is very different from (a) due to significant higher-order perturbations that are not involved in $Q_{\text{eff}}$. Fourier transformation is performed by using $\ell{=}10^{4}$ Floquet cycles. (c) Obtained from spin-dynamics simulation of Eq. (3) with $T{=}0.4$, $J_{2}{=}{-}1$, $h_{x}{=}1$, $h_{y}{=}1$ and $J_{x}{=}1$. The gapped mode still matches our theory, while the gapless mode disappears. Fourier transformation is performed by using $\ell{=}5\times 10^{4}$ Floquet cycles. The Fourier density of $q=0$ is illustrated in (d), where the gap is clearly visible. System size is $N=1000$. Ensemble averaging over $100$ realizations is performed with $\delta\theta_{j}{=}0.01\pi$ and $\delta\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$ for (a) and (b), while $\delta\phi_{j}$ randomly sampled within $(-\pi/6,\pi/6)$ for (c).
@@
Figure S12: (a) Linear dependence of $\omega$ versus $T$ for the gap size. Deviation between the numerical and theoretical results reduces for higher driving frequencies. (b) The width $\Gamma$ exhibits a power-law scaling in the high frequency regime with the scaling exponent $\alpha\approx 8$. We use $N=1000$, $J_{2}=-1$, $h_{x}=1$, $h_{y}=1$ and $J_{x}=1$. Ensemble averaging over $100$ realizations is performed with $\delta\theta_{j}{=}0.01\pi$ and $\delta\phi_{j}$ randomly sampled within the range $(-\pi/6,\pi/6)$.
In Fig. S12(b), we plot the linewidth $\Gamma$ for $q=0$ by performing Fourier transform to the dynamics of total magnetizations over time, for varying $T$. The linewidth is determined by averaging the full-width at $10\%$, $15\%$, and $20\%$ of the maximum intensity of the Fourier peak, with their standard deviation serving as the error bar. Interestingly, $\Gamma$ exhibits a power-law scaling $T^{\alpha}$ with a large scaling exponent $\alpha\approx 8$, suggesting that the lifetime of the mNG mode can be notably changed by slightly tuning the driving frequency.
@@
SM 4 Trotterized Simulation Protocol

### v4

@@ 6 Different Scaling Exponent of Linewidth
In the main text, Fig. 4, we find that the linewidth of the DSF scales as $T^{3}$, and we give one possible explanation using a hydrodynamic theory. One key assumption we made is that the attenuation $D_{0}$, induced by explicit SB, is dominated by $Q^{(2)}$. Together with the scaling of the gap size, one estimates the damping rate as $\Gamma\sim T^{3}$. However, this scaling may not be universal, and it can depend on the driving parameters and the randomness in the initial state ensemble.
@@
Figure S5: The width $\Gamma$ exhibits a power-law scaling in the high frequency regime; the dashed line with has fitted scaling $\alpha\approx 4$. We use $N{=}1000$, $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}-2$ for numerical simulations. Ensemble averaging over 100100100 realizations is performed with $\theta_{j}{=}0.1\pi$ and azmuthal angles $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@ 7 Thermalization and Hierarchical Symmetry Breaking in Prethermal Plateaus
Floquet systems generally heat up to infinite temperature, due to the absence of energy conservation. However, this can take a long time, with a timescale that normally increases exponentially with the driving frequency for systems with local interactions [65, 66]. In Fig. 1 in the main text, to show the persistent periodic oscillation of the magnetization we use a small $T$, which indeed significantly delays the onset of heating. During the time window that we can numerically simulate, the system does not heat up to infinite temperature, but stays in a prethermal plateau. Here, instead, we use a smaller driving frequency and stronger symmetry breaking strength $J_{x}$ to speed up heating. As shown in Fig. S6, after the initial transient oscillation, the system relaxes towards a prethermal plateau. Eventually, magnetization in three directions decays to zero, confirming the eventual heat death.
Figure S6: Dynaimics of the 1D driven Heisenberg-type model, Eq. (3), that shows the eventual thermalization. We use $N{=}1000$, $T{=}2.4$ $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}-2$, $J_{x}{=}5$ for numerical simulations. Ensemble averaging over 100100100 realizations is performed with $\theta_{j}{=}0.1\pi$ and azmuthal angles $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@
$U_{F}=\left(e^{-iH_{0}\frac{T}{6}}e^{-iH_{1}\frac{T}{6}}\right)e^{-iH_{2}\frac{T}{6}}\left(e^{iH_{0}\frac{T}{6}}e^{iH_{1}\frac{T}{6}}\right)e^{-iH_{2}\frac{T}{6}},$ (S.19)
@@
$H_{2}=-J_{2}\sum_{j}\bm{S}_{j}\cdot\bm{S}_{j+1},\quad\ H_{1}=-H_{2}+J_{1}\sum_{j\in even}S_{j}^{z}S_{j+2}^{z},\quad\ H_{0}=J_{x}\sum_{j\in odd}S_{j}^{x}S_{j+2}^{x}.$ (S.20)
[…]
Fig. S7 (a) shows the numerical time evolution of the total magnetization. From a sufficiently random initial ensemble, all components of the magnetization quickly settle into prethermal plateaus before eventually heating up to infinite temperature. A hierarchy in the prethermal lifetime is observed, with the $z$-component exhibiting a more stable plateau that persists longer than those of the $x$ and $y$ components. Moreover, Ref. [58] predicts that the prethermal lifetime $\tau$ can be parametrically prolonged by using smaller driving period $T$, $\tau_{x/y}\sim T^{-2}$ and $\tau_{z}\sim T^{-4}$, following a Fermi Golden Rule argument. There, this scaling has been numerically verified in quantum systems of a small size. In Fig. S7 (b) we also observe this scaling law in our classical systems where finite-size effects are negligible.
@@ 8 Classical System with Three-body Interactions
To realize emergent mNGs, explicit SB processes in $Q^{(1)}$ are not limited to being a generator of the non-Abelian group $G_{2}$, as long as $Q^{(1)}$ still preserves a subsymmetry of $G_{2}$. Here, we present an explicit example that realizes the symmetry-breaking process, $O(3)\to O(2)\to E$, and $Q^{(1)}$ involves three-body interactions which is not a simple generator of the $O(3)$ group. We still observe emergent mNGs with controllable dispersion relation that can be tuned by $T$. However, those mNGs have a shorter lifetime that scales linearly with the driving frequency.
@@
We obtain the effective Hamiltonian up to order $\mathcal{O}(T)$ as
@@
Following the definition of $Q^{(0)}$ and $Q^{(1)}$ in Eq. (2), $Q^{(0)}$ preserves $O(3)$ and $Q^{(1)}$ contains only the exchange terms and the $z$-component terms, therefore reducing the symmetry to $O(2)$ around the $z$-axis. We take $J_{2}>0$ such that the ground state features the FM order along the positive (negative) $z$-direction if $J_{1}$ and $h_{y}$ have the same (opposite) sign. The EOM generated by the truncated Hamiltonian Eq. (S.22) read as
@@
$\omega=\pm\left[\frac{2J_{2}}{5}(1-\cos{qa})+\frac{J_{1}h_{y}T}{200}(3+2\cos{qa}+2\cos{2qa})\right].$ (S.25)
A quadratic dispersion appears for modes of long wavelength when $q\rightarrow 0$ and a gap is now opened at $q=0$ of size $7T|J_{1}h_{y}|/200$ that is linear in $T$.
@@
Figure S8: Dynamics of the total magnetization in a 1D driven three-body interaction model, cf. Eq. (S.21). Magnetization in the $(x,y)$-plane (red and blue) rotates periodically, confirming the existence of mNGs. $\langle S_{z}^{\text{tot}}\rangle/N$ (green) oscillates on top of the conserved value (dashed line) due to higher order effects. We use a system of size $N=1000$ and parameters $T{=}1$, $J_{2}{=}1$, $h_{y}{=}1$, $J_{1}{=}1$, $J_{x}{=}1$. We perform ensemble averages over 100100100 realizations where initial spins deviate from the $z$-axis by one small angle $\theta_{j}=0.05\pi$, and their azimuthal angles $\phi_{j}$ are randomly sampled in the range $(-\pi/3,\pi/3)$.
@@
[…]
Figure S11: 1D driven three-body interaction model, cf. Eq. (S.21): (a) The linewidth $\Gamma$ exhibits a power-law scaling in the high-frequency regime with the exponent close to 111. The results of the quenched protocol closely mimic the Floquet protocol in the high-frequency regime. Floquet protocol generates a smaller linewidth for large $T$. We used $\theta_{j}{=}0.1\pi$. (b) The monotonic dependence of the ratio of linewidth to gap size, $\lambda$, on initial state parameter $\theta$. The ratio $\lambda$ can be suppressed by applying smaller spatial randomness. Various $\theta_{j}=\theta$ are used for characterizing the spatial randomness. We use $N{=}1000$, $T{=}1$, $J_{2}{=}1$, $h_{x}{=}1$, $J_{1}{=}1$ and $J_{x}{=}1$. 100100100 ensemble averages are performed with azimuthal angles $\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$.
@@ 9 Ferromagnetic Quantum Spin Model
[…]
@@
[…]
@@
We have provided a comprehensive analysis of FM in the main text, categorized as Case 1. For Case 2, we investigate the properties of the mNG modes in the Floquet system with antiferromagnetic (AFM) interaction, $J_{2}<0$. Significant distinctions between the two cases can be observed. For AFM, if we consider only the first two orders of the effective Hamiltonian, there exists a gapless Goldstone mode with linear dispersion and a gapped mNG with quadratic dispersion  [9]. Taking higher-order terms into account, the gapless mode disappears and the spectrum of the gapped mode is slightly modified. The gap linearly depends on the drive period $T$ in the high-frequency regime. Furthermore, we note that the stability of the gapped mode is quite different from the FM case.
@@ 1 Spin-wave Theory
[…]
@@ 2 Numerical Verification of Massive Goldstone Modes
Figure S13: Dynamical structure factor (DSF) through the Brillouin zone of AFM model. (a) Obtained from spin-dynamics simulation of a time-independent Hamiltonian, Eq. (S.27), with parameters $J{=}{-}1/5$, $h_{z}{=}1/25$. A gapped mode with quadratic dispersion and a gapless mode with linear dispersion appear in accordance with the theoretical prediction. (b) Floquet spin-dynamics simulation with $T{=}2$, $J_{2}{=}{-}1$, $h_{x}{=}2$, $h_{y}{=}2$ and $J_{x}{=}1$, which gives the same parameters as (a). However, this spectrum is very different from (a) due to significant higher-order perturbations that are not involved in $Q_{\text{eff}}$. Fourier transformation is performed by using $\ell{=}10^{4}$ Floquet cycles. (c) Obtained from spin-dynamics simulation of Eq. (3) with $T{=}0.4$, $J_{2}{=}{-}1$, $h_{x}{=}1$, $h_{y}{=}1$ and $J_{x}{=}1$. The gapped mode still matches our theory, while the gapless mode disappears. Fourier transformation is performed by using $\ell{=}5\times 10^{4}$ Floquet cycles. The Fourier density of $q=0$ is illustrated in (d), where the gap is clearly visible. System size is $N=1000$. Ensemble averaging over 100100100 realizations is performed with $\delta\theta_{j}{=}0.01\pi$ and $\delta\phi_{j}$ randomly sampled within $(-\pi/3,\pi/3)$ for (a) and (b), while $\delta\phi_{j}$ randomly sampled within $(-\pi/6,\pi/6)$ for (c).
@@
Figure S14: (a) Linear dependence of $\omega$ versus $T$ for the gap size. Deviation between the numerical and theoretical results reduces for higher driving frequencies. (b) The width $\Gamma$ exhibits a power-law scaling in the high frequency regime with the scaling exponent $\alpha\approx 8$. We use $N=1000$, $J_{2}=-1$, $h_{x}=1$, $h_{y}=1$ and $J_{x}=1$. Ensemble averaging over 100100100 realizations is performed with $\delta\theta_{j}{=}0.01\pi$ and $\delta\phi_{j}$ randomly sampled within the range $(-\pi/6,\pi/6)$.
[…]
@@
SM 4 Explicit Breaking of Abelian Symmetries
In the main text, we focus on the appearance of mNGs due to the breaking of non-abelian symmetries. In fact, mNGs can also exist in systems with an approximate abelian continuous symmetry. However, in this case, mass of mNGs becomes model-dependent and its determination can involve complicated calculations, both numerically and analytically.
For illustration, we consider the 2D antiferromagnetic transverse-field XXZ model on a square lattice, described by the Hamiltonian
$H=\sum_{\langle i,j\rangle}J_{\perp}(S_{i}^{x}S_{j}^{x}{+}S_{i}^{y}S_{j}^{y}){+}J_{z}S_{i}^{z}S_{j}^{z}{+}h\sum_{j}S_{j}^{x},$ (S.35)
with the $U(1)$ symmetry along $z$-direction. When $J_{z}/J_{\perp}<1$ and $h=0$, its ground state features easy-plane antiferromagnetic order, and there is a gapless NG mode with linear dispersion. A finite transverse field degrades the symmetry, $U(1)\to Z_{2}$, and a mNG mode appears with quadratic dispersion. The gap size can be analytically determined by linear spin-wave theory, exhibiting a complicated dependence on the parameters of the Hamiltonian, see details in Ref. [105]. One can further apply an additional field along the z-axis to break the remaining $Z_{2}$ symmetry, which simply induces an overall energy shift to the mNG spectrum.
In our work, we mainly focus on the non-abelian symmetry and the mass of mNGs is determined by the non-abelian group structure. Therefore, our Floquet protocol is generally applicable to different physical systems of interest and the prediction of the mass is less model-specific.
SM 5 Trotterized Simulation Protocol